Write all the factors of the following numbers: 
12 ????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

12 = 1 × 12 12 = 2 × 6 12 =3 ×4 12 = 4 x 3
Stop here since 3 and 4 have occurred earlier
Hence, the factors of 12 are 1, 2, 3, 4, 6 and 12

Thanks